OpenStudy (anonymous):

If a factory continuously dumps pollutants into a river at the rate of the quotient of the square root of t and 40 tons per day, then the amount dumped after 7 days is

OpenStudy (anonymous):

https://gyazo.com/e2dd9b1bf63a09b6cd39998bd19c7e47

OpenStudy (anonymous):

\[7*\frac{ \sqrt{7} }{ 40}\] @jim_thompson5910 is there even enough info to solve the problem?

OpenStudy (anonymous):

that 7 under the radical is supposed to be a t

OpenStudy (anonymous):

\[7*\frac{ \sqrt{t} }{ 40}\]

OpenStudy (anonymous):

Or would I take the integral of \[\frac{ \sqrt{t} }{ 40 }\] from 0 to 7?

jimthompson5910 (jim_thompson5910):

yes take the integral from 0 to 7

OpenStudy (anonymous):

0.309

jimthompson5910 (jim_thompson5910):

yep